# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

### v2

@@ 3.2 Existence result
theoremthmexistence Let Assumptions 1 and 2 hold and suppose $\mathcal{C}\not$$\emptyset$. Then there exists a constrained Nash equilibrium $x^{\star}\in\mathcal{X}$. The full proof of Theorem 3.2 is given in Appendix D; here we sketch the main elements.
A key challenge in our setting is the nonconvexity of the feasible region $\mathcal{C}$. Fixed point theorems such as Brouwer’s or Kakutani’s, which are typically used in existence proofs for Nash equilibria, require the feasible set to be convex. In order to establish the existence of a fixed point for a best response map defined over our nonconvex $\mathcal{C}$, we rely on a result that goes back to Eilenberg and Montgomery, (1946), and Begle, (1950), which relaxes the convexity requirement to the weaker topological notion of contractibility. A central part of our proof, Lemma 3.2, establishes that under Assumption 2, playerwise concavity of constraints implies contractibility of each connected component of $\mathcal{C}$.
Below, we define contractibility and state a version of Begle, (1950)’s theorem applied to finite-dimensional Euclidean space. For examples of contractible sets, and a proof of Fact 3.2, see Appendix C. {restatable}definitiondefcontractibility Let $S\subseteq\mathbb{R}^{n}$. We say $S$ is contractible if there exists $s_{0}\in S$, and a continuous map $H:[0,1]\times S\to S$ such that for all $s\in S$, it holds that $H(0,s)=s$ and $H(1,s)=s_{0}$. {restatable}factfactfp Let $X\subset\mathbb{R}^{n}$ be compact, contractible, and $\phi:X\to X$ continuous. Then there exists $x\in X$ such that $x=\phi(x)$, that is, $x$ is a fixed point of $\phi$.
As seen in Example 2, the feasible region $\mathcal{C}$ may be disconnected; thus, it cannot be contractible. However, we establish the following key contractibility result.
@@
lemmalemcontract Suppose Assumptions 1 and 2 hold. Let $\mathcal{C}\not$$\emptyset$ and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$. Then $\mathcal{C}^{\prime}$ is contractible.
@@
Property (a) is shown via convexity of the slices $\mathcal{C}^{\prime}_{i}(x_{-i})$ which follows directly from the playerwise concavity of the constraints. Regarding (b), we observe that $\overline{c}(\cdot)$ and $\mathcal{C}_{i}^{\prime}(\cdot)$, from which we compose $H$, are both continuous. In particular, continuity of the set-valued mapping $\mathcal{C}_{i}^{\prime}(\cdot)$ follows from standard results in parametric optimization, which apply under the MFCQ condition (Assumption 2). For (c), by construction, for all $x\in\mathcal{C}^{\prime}$, $H(0,x)=x$ and $H(1,x)=\overline{c}(\operatorname{supp}(\mathcal{C}_{1}^{\prime}))$.
@@ 4.1 Constrained potential games
While Assumptions 1 and 2 suffice for existence, without additional assumptions on the game structure, the problem of finding $\epsilon$-approximate Nash equilibria is known to be PPAD-hard even for unconstrained normal-form games (Daskalakis et al.,, 2009; Rubinstein,, 2015). Therefore, we focus on the subclass of potential games (Monderer and Shapley,, 1996) with coupling constraints. For our purpose, it suffices to have a potential function defined over the feasible region.
Definition 1.
@@
Examples of such continuous potential games trivially include cooperative games, and, as shown in Example 3 below, mixed strategy extensions of finite action potential games with coupling constraints. This class of potential games also arises in economics, e.g. in the context of the Cournot competition (Cournot,, 1838), in communication networks (Scutari et al.,, 2006), and in transportation (Altman and Wynter,, 2004).
Example 3.
@@ 4.3 Interior point method
Access to a feasible initial strategy is crucial, as otherwise even a two-player game with bilinear constraints requires solving an NP-hard optimization problem (Witsenhausen,, 1986). For iterate $x^{(t)}\in\mathcal{C}$, we denote its feasibility margin by $\beta^{(t)}\coloneqq\min_{j\in[b]}\{c_{j}(x^{(t)})-\alpha_{j}\}$.
Assumption 3.
@@
To ensure feasibility of the iterates during the algorithm, we use log barrier regularized gradient descent, an interior point approach recently studied for nonconvex problems with nonconvex constraints (Hinder and Ye,, 2024; Usmanova et al.,, 2024). Under the assumed potential structure, we show that independently updated iterates remain in the interior of the joint feasible region.
@@ 4.4 Convergence Result
In order to prove our convergence result, we adopt standard assumptions on smoothness and Lipschitz continuity of utility functions, see also Mertikopoulos et al., (2024), and extend these regularity conditions to the constraints.
Assumption 4.
@@
where $M^{\eta,\beta^{(t)}}_{\Phi}\coloneqq M\sqrt{m}+\eta Mb(\beta^{(t)})^{-1}+\eta L^{2}b(\beta^{(t)})^{-2}$. The first term in (2) ensures that iterates remain within the feasible region. The second term bounds $\gamma^{(t)}$ by the inverse of the local smoothness parameter $M^{\eta,\beta^{(t)}}_{\Phi}$. We further describe these terms in the proof overview of Theorem 2 below.
While for the existence result, the MFCQ condition stated in Assumption 2 is sufficient, our convergence proof requires an extended version of this condition. Namely, a uniform constraint ascent direction must exist for all feasible joint strategies $x\in\mathcal{C}$ near the boundary of $\mathcal{C}$, ensuring that log barrier gradients consistently steer the iterates away from the boundary.
Assumption 5.
@@
Note that for $\rho=0$ and $\mathcal{X}=\mathbb{R}^{n}$, we recover the classic Mangasarian-Fromovitz constraint qualification (MFCQ) condition (Mangasarian and Fromovitz,, 1967).
@@
Now, we are ready to state our main theorem about convergence to a constrained Nash equilibrium. {restatable}theoremthmpotential Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 1, 3, 4, and 5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (2), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium. We first discuss how this result relates to the existing literature. Then we provide an overview of our main ideas and the challenges of our analysis; for the full proof we refer to Appendix F.
Remarks.
@@
    The work of Anagnostides et al., (2022) also considers concave continuous action potential games, but without coupling constraints, establishing convergence to a Nash equilibrium at an $\mathcal{O}(\epsilon^{-2})$ best-iterate rate (see their Theorem B.6). Thus, handling coupling constraints via log barrier regularization comes at the cost of an additional $\mathcal{O}(\epsilon^{-1})$ factor. This is due to small adaptive stepsizes as iterates approach the boundary of the feasible region.
@@
    While we are not aware of any results on learning Nash equilibria in continuous action potential games with coupling constraints, a related result exists for constrained Markov potential games under an independent learning protocol similar to ours (Jordan et al.,, 2024). Unlike our approach, this result relies on a double-loop proximal-point-like algorithm that requires a so-called uniform Slater’s condition. Intermediate iterates are not guaranteed to remain within the feasible region. In the exact-gradient setting, a best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium at a slower $\mathcal{O}(\epsilon^{-4})$ rate is proven.
@@
Next, we reason about feasibility of the joint strategy iterates and smoothness of $\Phi^{\eta}$ along the gradient ascent trajectory. For $\beta^{(t)}$-feasible $x^{(t)}$, we give a Hessian bound $\lVert\nabla^{2}\Phi^{\eta}(x)\rVert\leq M^{\eta,\beta^{(t)}}_{\Phi}$. Moreover, we show that if an iterate $x^{(t)}$ is $\beta^{(t)}$-feasible, and stepsizes are chosen as in (2), then all strategies on the line segment $x^{(t)},x^{(t+1)}]$ are $(\beta^{(t)}/2)$-feasible. Consequently, $\Phi^{\eta}$ is $(M_{\Phi}^{\eta,\beta^{(t)}/2})$-smooth over $x^{(t)},x^{(t+1)}]$. Therefore, if we additionally ensure $\gamma^{(t)}\leq 1/(M_{\Phi}^{\eta,\beta^{(t)}/2})$, as in (2), a standard inequality from nonconvex smooth optimization (Bubeck et al., (2015), Lemma 3.6) shows sufficient increase proportional to $\gamma^{(t)}$. Nevertheless, $M_{\Phi}^{\eta,\beta^{(t)}/2}$ may grow arbitrarily and lead to diminishing stepsizes. Leveraging Assumption 5, we give a lower bound on $\gamma^{(t)}$ based on the fact that the log barrier term dominates and steers iterates away from the boundary of $\mathcal{C}$ whenever the feasibility margin $\beta^{(t)}$ is too small.
@@ 5 SIMULATIONS
To demonstrate the convergence of our method, we conduct numerical simulations on two continuous potential games with shared coupling constraints. The first is a simple two-player cooperative game played over a nonconvex feasible region. The second is a routing problem with link capacity constraints. We measure convergence in terms of $\text{Nash-Gap}_{i}(x)\coloneqq\max_{x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})}u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x)$ for each player $i$. This can be computed by solving the single-player constrained optimization problem where $x_{-i}$ remains fixed. We further define $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)$. This serves as a measure of convergence, as for any $\epsilon>0$, if $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)\leq\epsilon$, then $x$ is a constrained $\epsilon$-approximate Nash equilibrium.
@@
Figure 4: Nash gaps and constraint values over iterations $t$ for the cooperative game.
@@
Beyond the cooperative case, we turn to a classical distributed routing problem, a canonical example of a potential game well studied in the literature on network routing (Menache and Ozdaglar,, 2011). We extend this model by incorporating coupled capacity constraints. Specifically, 5 players, one corresponding to each color in Figure 5, route some amount of flow through 5 paths in the network, each associated to the respective player ($R_{1}$: blue, $R_{2}$: red, $R_{3}$: brown, $R_{4}$: orange, $R_{5}$: green). For $i\in[5]$, we let $x_{i}\in[0,10]$ denote the flow on player $i$’s path $R_{i}$.
Figure 5: Routing game (Tatarenko, (2017), Ex. 2.2).
@@
with upper bounds $\alpha_{1}=2.7$ and $\alpha_{2}=7$. Note that by negating, these constraints can be formulated as lower bounds in order to match our problem definition. Moreover, utilities and constraints are playerwise concave due to linearity and convexity of the functions $P_{l}$ for $l\in[5]$, hence satisfying Assumption 1.
@@
Figure 6: Nash gap and constraint values over iterations $t$ for the routing game.
@@ 6 CONCLUSION
Our findings suggest several directions for future work. Can equilibrium existence and computation be characterized when utilities or constraints lack playerwise concavity, as in Markov games? Does our algorithmic approach extend to settings with stochastic zeroth- or first-order feedback? And finally, we ask whether our approach can be applied to learning equilibria in other tractable subclasses of playerwise concave games, such as zero-sum or monotone games.
Supplementary Materials
@@ Appendix B Discussion on Constraint Structure
We point out that in above definition, deviations of player $i$ only need to be feasible with respect to player $i$’s constraints; this is what distinguishes non-shared constraints from shared ones. Note that this equilibrium notion matches the one introduced by Debreu, (1952). The nonemptiness assumption of Debreu, (1952) then requires that for all $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, we have $\mathcal{C}_{i}(x_{-i})\not$$\emptyset$. Under this assumption, Debreu, (1952) proves existence of constrained Nash equilibria.
Is it possible to extend Debreu, (1952)’s theorem and establish existence results based on playerwise concavity (our Assumption 1) in this class of games with non-shared constraints, without introducing further assumptions such as the above nonemptiness condition on the feasible response sets? We give a negative answer in the form of the following counterexample. Note that the counterexample even has jointly linear constraints and utilities.
Proposition 1.
@@
with thresholds $\alpha_{1,1}=\alpha_{1,2}=0.5$. Assumption 1 on the playerwise concavity of utilities and constraints holds due to linearity of utility and constraint functions. The joint feasible region is
@@
Note that Debreu, (1952)’s assumption prevents constructing such pathological cases: In above example, the nonemptiness condition is violated, as for any $\hat{x}\in[0,0.5)^{2}$, $\widetilde{\mathcal{C}}_{1}(\hat{x}_{2})=\widetilde{\mathcal{C}}_{2}(\hat{x}_{1})=\emptyset$. However, as seen in our Example 2, Debreu, (1952)’s nonemptiness assumption is strong and captures only a limited class of coupling constraints.

### v3

@@ 3.2 Existence Result
theoremthmexistence Let Assumptions 2.2 and 3.1 hold, let $\mathcal{C}\not$$\emptyset$. Then there exists a constrained Nash equilibrium $x^{\star}\in\mathcal{X}$. The full proof of Theorem 3.2 is given in Appendix D; here we sketch the main elements.
A key challenge in our setting is the nonconvexity of the feasible region $\mathcal{C}$. Fixed point theorems such as Brouwer’s or Kakutani’s, which are typically used in existence proofs for Nash equilibria, require the feasible set to be convex. In order to establish the existence of a fixed point for a best response map defined over our nonconvex $\mathcal{C}$, we rely on a result that goes back to Eilenberg & Montgomery (1946), and Begle (1950), which relaxes the convexity requirement to the weaker topological notion of contractibility. A central part of our proof, Lemma 3.2, establishes that under Assumption 3.1, playerwise concavity of constraints implies contractibility of each connected component of $\mathcal{C}$.
Below, we define contractibility and state a version of Begle (1950)’s theorem applied to finite-dimensional Euclidean space. For examples of contractible sets, and a proof of Fact 3.2, see Appendix C. {restatable}definitiondefcontractibility Let $S\subseteq\mathbb{R}^{n}$. We say $S$ is contractible if there exists $s_{0}\in S$, and a continuous map $H:[0,1]\times S\to S$ such that for all $s\in S$, it holds that $H(0,s)=s$ and that $H(1,s)=s_{0}$. {restatable}factfactfp Let $X\subset\mathbb{R}^{n}$ be compact, contractible, and let $\phi:X\to X$ be continuous. Then there exists $x\in X$ such that $x=\phi(x)$, that is, $x$ is a fixed point of $\phi$.
As seen in Example 3.3, the feasible region $\mathcal{C}$ may be disconnected; thus, it cannot be contractible. However, we establish the following key contractibility result.
@@
lemmalemcontract Suppose Assumptions 2.2 and 3.1 hold. Let $\mathcal{C}\not$$\emptyset$ and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$. Then $\mathcal{C}^{\prime}$ is contractible.
@@
Property (a) is shown via convexity of the slices $\mathcal{C}^{\prime}_{i}(x_{-i})$ which follows directly from the playerwise concavity of the constraints. Regarding (b), we observe that $\overline{c}(\cdot)$ and $\mathcal{C}_{i}^{\prime}(\cdot)$, from which we compose $H$, are both continuous. In particular, continuity of the set-valued mapping $\mathcal{C}_{i}^{\prime}(\cdot)$ follows from standard results in parametric optimization, which apply under the MFCQ condition (Assumption 3.1). For (c), by construction, for all $x\in\mathcal{C}^{\prime}$, $H(0,x)=x$ and $H(1,x)=\overline{c}(\operatorname{supp}(\mathcal{C}_{1}^{\prime}))$.
@@ 4.1 Constrained Potential Games
While Assumptions 2.2 and 3.1 suffice for existence, without additional assumptions on the game structure, the problem of finding $\epsilon$-approximate Nash equilibria is known to be PPAD-hard even for unconstrained normal-form games (Daskalakis et al., 2009; Rubinstein, 2015). Therefore, we focus on the subclass of potential games (Monderer & Shapley, 1996) with coupling constraints. For our purpose, it suffices to have a potential function defined over the feasible region.
Definition 4.1.
@@
Examples of such continuous potential games trivially include cooperative games, and, as shown in Example 4.2 below, mixed strategy extensions of finite action potential games with coupling constraints. This class of potential games also arises in economics, e.g. in the context of the Cournot competition (Cournot, 1838), in communication networks (Scutari et al., 2006), and in transportation (Altman & Wynter, 2004).
Example 4.2.
@@ 4.3 Interior Point Method
Access to a feasible initial strategy is crucial, as otherwise even a two-player game with bilinear constraints requires solving an NP-hard optimization problem (Witsenhausen, 1986). For iterate $x^{(t)}\in\mathcal{C}$, we denote its feasibility margin by $\beta^{(t)}\coloneqq\min_{j\in[b]}\{c_{j}(x^{(t)})-\alpha_{j}\}$.
Assumption 4.3.
@@
To ensure feasibility of the iterates during the algorithm, we use log barrier regularized gradient descent, an interior point approach recently studied for nonconvex problems with nonconvex constraints (Hinder & Ye, 2024; Usmanova et al., 2024). Under the assumed potential structure, we show that independently updated iterates remain in the interior of the joint feasible region.
@@ 4.4 Convergence Result
In order to prove our convergence result, we adopt standard assumptions on smoothness and Lipschitz continuity of utility functions, see also Mertikopoulos et al. (2024), and extend these regularity conditions to the constraints.
Assumption 4.4.
@@
where $M^{\eta,\beta^{(t)}}_{\Phi}\coloneqq M\sqrt{m}+\eta Mb(\beta^{(t)})^{-1}+\eta L^{2}b(\beta^{(t)})^{-2}$. The first term in (2) ensures that iterates remain within the feasible region. The second term bounds $\gamma^{(t)}$ by the inverse of the local smoothness parameter $M^{\eta,\beta^{(t)}}_{\Phi}$. We further describe these terms in the proof overview of Theorem 4.6 below.
While for the existence result, the MFCQ condition stated in Assumption 3.1 is sufficient, our convergence proof requires an extended version of this condition. Namely, a uniform constraint ascent direction must exist for all feasible joint strategies $x\in\mathcal{C}$ near the boundary of $\mathcal{C}$, ensuring that log barrier gradients consistently steer the iterates away from the boundary.
Assumption 4.5.
@@
Note that for $\rho=0$ and $\mathcal{X}=\mathbb{R}^{n}$, we recover the classic Mangasarian-Fromovitz constraint qualification (MFCQ) condition (Mangasarian & Fromovitz, 1967).
@@
Now, we are ready to state our main theorem about convergence to a constrained Nash equilibrium. {restatable}theoremthmpotential Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 2.2, 4.3, 4.4, and 4.5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (2), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium. We first discuss how this result relates to the existing literature. Then we provide an overview of our main ideas and the challenges of our analysis; for the full proof we refer to Appendix F.
Remarks
@@
    The work of Anagnostides et al. (2022) also considers concave continuous action potential games, but without coupling constraints, establishing convergence to a Nash equilibrium at an $\mathcal{O}(\epsilon^{-2})$ best-iterate rate (see their Theorem B.6). Thus, handling coupling constraints via log barrier regularization comes at the cost of an additional $\mathcal{O}(\epsilon^{-1})$ factor. This is due to small adaptive stepsizes as iterates approach the boundary of the feasible region.
@@
    While we are not aware of any results on learning Nash equilibria in continuous action potential games with coupling constraints, a result related to ours exists for constrained Markov potential games under an independent learning protocol (Jordan et al., 2024). Unlike our approach, this result relies on a double-loop proximal-point-like algorithm that requires a so-called uniform Slater’s condition. Intermediate iterates are not guaranteed to remain within the feasible region. In the exact-gradient setting, a best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium at a slower $\mathcal{O}(\epsilon^{-4})$ rate is proven.
@@
Next, we reason about feasibility of the joint strategy iterates and smoothness of $\Phi^{\eta}$ along the gradient ascent trajectory. For $\beta^{(t)}$-feasible $x^{(t)}$, we give a Hessian bound $\lVert\nabla^{2}\Phi^{\eta}(x)\rVert\leq M^{\eta,\beta^{(t)}}_{\Phi}$. Moreover, we show that if an iterate $x^{(t)}$ is $\beta^{(t)}$-feasible, and stepsizes are chosen as in (2), then all strategies on the line segment $x^{(t)},x^{(t+1)}]$ are $(\beta^{(t)}/2)$-feasible. Consequently, $\Phi^{\eta}$ is $(M_{\Phi}^{\eta,\beta^{(t)}/2})$-smooth over $x^{(t)},x^{(t+1)}]$. Therefore, if we additionally ensure $\gamma^{(t)}\leq 1/(M_{\Phi}^{\eta,\beta^{(t)}/2})$, as in (2), a standard inequality from nonconvex smooth optimization (Bubeck et al. (2015), Lemma 3.6) shows sufficient increase proportional to $\gamma^{(t)}$. Nevertheless, $M_{\Phi}^{\eta,\beta^{(t)}/2}$ may grow arbitrarily and lead to diminishing stepsizes. Leveraging Assumption 4.5, we give a lower bound on $\gamma^{(t)}$ based on the fact that the log barrier term dominates and steers iterates away from the boundary of $\mathcal{C}$ whenever the feasibility margin $\beta^{(t)}$ is too small.
@@ 5 Simulations
To demonstrate the convergence of our method, we conduct numerical simulations on two continuous potential games with shared coupling constraints. The first is a simple two-player cooperative game played over a nonconvex feasible region. The second is a routing problem with link capacity constraints. We measure convergence in terms of $\text{Nash-Gap}_{i}(x)\coloneqq\max_{x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})}u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x)$ for each player $i$. This can be computed by solving the single-player constrained optimization problem where $x_{-i}$ remains fixed. We define $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)$. This serves as a measure of convergence, as for any $\epsilon>0$, if $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)\leq\epsilon$, then $x$ is a constrained $\epsilon$-approximate Nash equilibrium.
@@
Figure 4: Convergence of log barrier regularized gradient ascent in the cooperative game: individual Nash gaps $\smash{\text{Nash-Gap}_{i}(x^{(t)})}$ decrease to zero while the constraint $\smash{c(x^{(t)})}$ remains below the threshold $\alpha$, indicating feasibility throughout.
@@
Beyond the cooperative case, we turn to a classical distributed routing problem, a canonical example of a potential game well studied in the literature on network routing (Menache & Ozdaglar, 2011). We extend this model by incorporating coupled capacity constraints. Specifically, we adopt an example of Tatarenko (2017); see their Example 2.2. Five players, one corresponding to each color in Figure 5, aim to route some amount of flow through five paths in the network, each associated to the respective player ($R_{1}$: blue, $R_{2}$: red, $R_{3}$: brown, $R_{4}$: orange, $R_{5}$: green). For $i\in[5]$, we let $x_{i}\in[0,10]$ denote the flow on player $i$’s path $R_{i}$.
Figure 5: Routing game (Tatarenko (2017), Ex. 2.2).
@@
with upper bounds $\alpha_{1}=2.7$ and $\alpha_{2}=7$. Note that by negating, these constraints can be formulated as lower bounds in order to match our problem definition. Moreover, utilities and constraints are playerwise concave due to linearity and convexity of the functions $P_{l}$ for $l\in[5]$, hence satisfying Assumption 2.2.
@@
Figure 6: Convergence in the constrained routing game: the Nash gap decreases to zero while link capacity constraints $c_{1}$ and $c_{2}$ remain below their respective thresholds $\alpha_{1}$ and $\alpha_{2}$ throughout all iterations.
@@ 6 Conclusion
Our findings suggest several directions for future work. First, characterizing equilibrium existence and computation in settings where utilities or constraints lack playerwise concavity, such as Markov games, remains to be further explored. Second, extending our algorithmic approach to accommodate stochastic zeroth- or first-order feedback is a natural next step. Third, adapting our framework to learn equilibria in other tractable subclasses of playerwise concave games, including zero-sum and monotone games, is another direction worth exploring.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Discussion on Constraint Structure
We point out that in above definition, deviations of player $i$ only need to be feasible with respect to player $i$’s constraints; this is what distinguishes non-shared constraints from shared ones. Note that this equilibrium notion matches the one introduced by Debreu (1952). The nonemptiness assumption of Debreu (1952) then requires that for all $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, we have $\mathcal{C}_{i}(x_{-i})\not$$\emptyset$. Under this assumption, Debreu (1952) proves existence of constrained Nash equilibria.
Is it possible to extend Debreu (1952)’s theorem and establish existence results based on playerwise concavity (our Assumption 2.2) in this class of games with non-shared constraints, without introducing further assumptions such as the above nonemptiness condition on the feasible response sets? We give a negative answer in the form of the following counterexample. Note that the counterexample even has jointly linear constraints and utilities.
Proposition B.1.
@@
with thresholds $\alpha_{1,1}=\alpha_{1,2}=0.5$. Assumption 2.2 on the playerwise concavity of utilities and constraints holds due to linearity of utility and constraint functions. The joint feasible region is
@@
Note that Debreu (1952)’s assumption prevents constructing such pathological cases: In above example, the nonemptiness condition is violated, as for any $\hat{x}\in[0,0.5)^{2}$, $\widetilde{\mathcal{C}}_{1}(\hat{x}_{2})=\widetilde{\mathcal{C}}_{2}(\hat{x}_{1})=\emptyset$. However, as seen in our Example 3.3, Debreu (1952)’s nonemptiness assumption is strong and captures only a limited class of coupling constraints.